Zinc metal reacts with hydrochloric acid produce zinc chloride and and hydrogen. When 0.103g of zinc is combined with enough hydrochloric acid to make 50mL of solution in a coffee cup calorimeter raising the temperature from 24.5 to 32.33 degrees celsius. Please calculate the heat flow and change in enthalpy.
Answer #1

Given

Mass of zinc combined = 0.103 g

Volume = 50 mL

Initial temperature = 24.50C

Final temperature = 32.330C

Now

Density of water = 1 g/mL

Density = Mass / Volume

==>

Mass = Density * Volume

= 1 g/mL * 50mL

= 50 g

Heat flow(Q) = m c ΔT

M is the mass

C is the specific heat

ΔT is change in temperature

So

M = 50g

C = 4.184 J/g0C

ΔT = (32.33-24.5) 0C

= 7.830C

Q = 50 g * 4.184 J/g0C * 7.830C

= 1638.036 J

We know that

Change in enthalpy (ΔH) = Q/n

Here

n is the no of moles

moles of zinc = mass of zinc/ molar mass of zinc

= 0.103 g/65.38 g/mol

= 1.575*10-3 moles

So

Change in enthalpy (ΔH) = 1638.036 J / 1.575*10-3 moles

= 1.04 * 106 J/mol